OpenStudy (anonymous):

find the quotient 3m2−27n2/3m+9n

OpenStudy (anonymous):

can someone help ill give you a medel

OpenStudy (campbell_st):

factor the numerator and denominator here is the 1st step \[\frac{3(m^2 - 9n^2)}{3(m + 3n)}\] remove the common factor... next the numerator is the difference of 2 squares and can be factored a^2 - b^2 = (a - b)(a + b) then remove the common factor you identify

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well when you cancel the 3 its \[\frac{m^2 - 9n^2}{m + 3n}\] factor the quadratic in the numerator

OpenStudy (anonymous):

i did all that but my answer was wrong @campbell_st

OpenStudy (campbell_st):

so what's you answer...?

OpenStudy (anonymous):

it was 7a +8b @campbell_st

OpenStudy (campbell_st):

wow.... thats funny given this question is m and n

OpenStudy (anonymous):

7m+8n i was typing to fast

OpenStudy (campbell_st):

ok... so here is a question... can you factor a^2 - b^2

OpenStudy (anonymous):

no i cant its new and everytime i tried to learn this we end up leaving class @campbell_st

OpenStudy (campbell_st):

ok... this quadratic is the difference of 2 squares... and can be factored as a^2 - b^2 = (a - b)(a + b) does that make sense..?

OpenStudy (anonymous):

3-27 then 3+9 @campbell_st

OpenStudy (campbell_st):

nope... you are factoring m^2 - 9n^2

OpenStudy (campbell_st):

\[m^2 - 9n^2 = (m - ?)(m + ?)\] what's the value of ? so that ? x ? = 9n^2

OpenStudy (anonymous):

i just need the answer to past @campbell_st

OpenStudy (anonymous):

my grade

OpenStudy (campbell_st):

whats the square root of 9...?

OpenStudy (anonymous):

3

OpenStudy (campbell_st):

yes... so you problem is \[\frac{3m^2 - 27n^2}{3m + 9m} = \frac{3(m - 3n)(m + 3n)}{3(m + 3n)}\] cancel the common factors and you'll have your answer

OpenStudy (anonymous):

i dont know how @campbell_st

OpenStudy (campbell_st):

outside the brackets... whats the same in the top and the bottom...?

OpenStudy (anonymous):

3mn

OpenStudy (campbell_st):